OpenStudy (anonymous):

Multiply 8 times the square root of 3 by 9 times the square root of 27

OpenStudy (solomonzelman):

\[8\sqrt{3} \times 9\sqrt{27}\] lets put numbers together and roots together. \[(8 \times 9 )\times( \sqrt{3} \times \sqrt{27})\]\[72 \times \sqrt{81}\]\[72 \times 9\]\[648\]